Stock split versus stock dividendFirm

Mammoth Corporation is considering a? 3-for-2 stock split. It currently has the? stockholders' equity position shown. The current stock price is? $120 per share. The most recent? period's earnings available for common stock is included in retained earnings.

Preferred stock Common stock (150,000 shares at $5 par) Paid-in capital in excess of par Retained earnings $1,000,000 750,000 1,700,000 10,000,000 $13,450,000 Total stockholders equity

a. What effect on? Mammoth's equity account would result from the stock split?

b. What change in stock price would you expect to result from the stock? split?

c. What is the maximum cash dividend per share that the firm could pay on common stock before and after the stock? split? ? (Assume that legal capital includes all?paid-in capital.)

d. Contrast your answers to parts a through c. with the circumstances surrounding a? 50% stock dividend.

e. Explain the differences between stock splits and stock dividends.
Answer #1

a. Before stock split, there are 150,000 stocks outstanding at $5 par value each. After 3-for-2 split:

total number of outstanding stocks = 150,000 * 3/2 = 225,000 (thus for every 2 stocks before, each inverstor will have 3)

With stock split, total par value will remain unchanged before and after, so par value per stock would be adjusted.

par value after stock split = $5 * 2/3 = 3.3333.

So total par value before and after split will be same equal to 750,000.

b. Before stock split, stock price was $120. Before and after stock split, market capitalization is expected to remain same to $120 * 150,000 = 18,000,000.

Stock Price after split = 18,000,000/225,000 = $80 (thus stock price after split decreases, becomes 2/3 rd of original)

Note: total market capitalization is still same =  $80 * 225,000 = 18,000,000.

c. Before Stock split, max dividend per share = Retained Earnings / 150,000 = 10,000,000 / 150,000 = $66.67

{considering retained earnings are undistributed reserve}

After Stock split, max dividend per share = Retained Earnings / 150,000 = 10,000,000 / 225,000 = $44.44

d. After 50% stock dividend:

total outstanding shares = 150,000 * 1.5 = 225,000

par value per share would remain same after stock dividend to be $5.

total par value would be = $5 * 225,000 = 1,125,000 (would increase)

{in case of stock dividend the total equity will remain same, but there will be some journal entries made to transfer some part of retained earning to Paid in capital excess of par & common stock account}

Stock price would be = 18,000,000/225,000 = $80 {so market capitalization is still same}

Retained earnings would decrease after stock dividend to (funds from retained earnings will be transferred to common stock account & Paid in capital excess of par accounts):

common stock increase = increase in total par value = par value after stock dividend - par value before stock dividend = 1,125,000 - 750,000 = 375,000 ...(1)

Paid in capital excess of par accounts increase = 120 * 75,000 - 375,000 = 8,625,000 ...(2) {incremental value - incremental par value}

Retained earnings after stock split = Retained earnings before stock dividend - common stock increase - Paid in capital excess of par increase = 10,000,000 - 375,000 - 8,625,000 = 1,000,000

max cash dividend per share after stock dividend = 1,000,000 / 225,000 = 4.4444

{So we note from above analysis, that stock split & stock dividend will result in same wealth to shareholder. Par value per share will decrease in case of stock split but not in case of stock dividend. In case of stock dividend there will be transfer of fund from retained earnings to common stock and excess paid in capital account. Total equity will however be same in both cases}